5) Four new designs for tooling re proposed. The organizations target rate on investment is 15%. No residual salvage value would exist for any of the alternatives. They would each have six years life. A B C D First cost 60,,000 80,000 100,000 40,000 Saving/year 15,000 25,000 31,000 12,000 Determine each alternative’s IRR and calculate incremental IRR (8 points)
Answer #1

Answer:

Calculation of NPV @ 10% and 30%

Particulars A B C D
First Cost 60000 80000 100000 40000
Cash outflow 60000 80000 100000 40000
Annual Saving 15000 25000 31000 12000
P/A 6 years 10% 4.355 4.355 4.355 4.355
Present Value 65325 108875 135005 52260
NPV @ 10% 5325 28875 35005 12260
P/A 6 years 20% 2.643 2.643 2.643 2.643
Present Value 39645 66075 81933 31716
NPV @ 30% -20355 -13925 -18067 -8284

IRR of A = Lower Rate + ( Positive NPV / Positive NPV - Negative NPV ) * ( Higher rate - Lower rate )

= 10 + ( 5325 / 5325 + 20355) * ( 30 - 10 )

= 10 + 0.21 ( 20 )

= 10 + 4.14

= 14.14 % is IRR of A

IRR of B = Lower Rate + ( Positive NPV / Positive NPV - Negative NPV ) * ( Higher rate - Lower rate )

= 10 + ( 28875 / 28875 + 13925 ) * ( 30 -10 )

= 10 + 0.67 ( 20 )

= 10 + 13.49

= 23.49 % is IRR of B

IRR of C = Lower Rate + ( Positive NPV / Positive NPV - Negative NPV ) * ( Higher rate - Lower rate )

= 10 + ( 35005 / 35005 + 18067 ) * ( 30 - 10 )

= 10 + 0.65 ( 20 )

= 10 + 13.19

= 23.19 % is IRR of C

IRR of D = Lower Rate + ( Positive NPV / Positive NPV - Negative NPV ) * ( Higher rate - Lower rate )

= 10 +(12260 / 12260 + 8284 ) * ( 30 - 10 )

= 10 + 0.60 ( 20 )

= 10 + 12

= 22 % is IRR of D